OpenStudy (anonymous):

Solve the equation. Log(x+7)- logs=3 ~3.5 ~142.7 ~0.0070 ~0.0707

OpenStudy (anonymous):

*-logx=3

Nnesha (nnesha):

log rules quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ---to-> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

Which one would I use?

Nnesha (nnesha):

look at the question which one would you use quotient rule or product ? (focus on sign) and there is no exponent so you can't apply power rule